# Polytope Volume Monitoring Problem: Formulation and Solution via Parametric Linear Program Based Control Barrier Function

### v1

@@ Abstract
Motivated by the latest research on feasible space monitoring of multiple control barrier functions (CBFs) as well as polytopic collision avoidance, this paper studies the Polytope Volume Monitoring (PVM) problem, whose goal is to design a control law for inputs of nonlinear systems to prevent the volume of some state-dependent polytope from decreasing to zero. Recent studies have explored the idea of applying Chebyshev ball method in optimization theory to solve the case study of PVM; however, the underlying difficulties caused by nonsmoothness have not been addressed. This paper continues the study on this topic, where our main contribution is to establish the relationship between nonsmooth CBF and parametric optimization theory through directional derivatives for the first time, so as to solve PVM problems more conveniently. In detail, inspired by Chebyshev ball approach, a parametric linear program (PLP) based nonsmooth barrier function candidate is established for PVM, and then, sufficient conditions for it to be a nonsmooth CBF are proposed, based on which a quadratic program (QP) based safety filter with guaranteed feasibility is proposed to address PVM problems. Finally, a numerical simulation example is given to show the efficiency of the proposed safety filter.
@@ I Introduction
The appearance of multiple CBFs and input constraints would impose great challenges in the feasibility of the resulting QP-based control laws, which has been paid much attention recently [5, 6, 7, 8, 9]. Some primary works study sufficient conditions of the compatibility of multiple CBFs [6, 7]. More recently, in [8], the nonsmooth CBF, dating backing to [5], is introduced as a systematic tool to address challenges in guaranteeing the feasibility and continuity. Meanwhile, a simpler way of addressing the feasibility problem appears in [9]. In detail, “Feasible CBF Space” is defined to describe the state-dependant polytope (in input space) formulated by CBF conditions and input constraints. Then, Feasible Space Monitoring (FSM) problem is established to monitor the feasibility of the resulting QP-based control laws, for which Chebyshev ball and inscribing ellipsoid methods are proposed to build parametric linear program (PLP) based CBFs. For engineering implementations, the differentiable optimization toolbox is applied to approximate the gradient of the resulting (nonsmooth implicit) barrier function candidate by some heuristic gradient values. As concluded in [9], the nonsmooth control theoretic considerations are left for future work. To the best of our knowledge, this problem has been open until now. Such observation motivates this paper.
@@
Meanwhile, researchers have been gradually interested in building CBFs via parametric optimization-defined metrics for related problems, such as network connectivity maintenance [10] and polytopic collision avoidance (PCA) [11, 12, 13, 14, 15, 16]. However, it is difficult to extend these results to solve the above open problem in FSM. Roughly speaking, there are three potential ways, yet all meet significant difficulties: 1) In [10], the Laplacian’s eigenvalue is chosen as a barrier function candidate, then, a nonsmooth CBF condition in the generalized gradient form is formulated. The idea of building nonsmooth CBF conditions based on generalized gradient can only be applied to specific problems, and it is not applicable to other parametric optimization problems. 2) In [15, 16], the minimum distance field for PCA, i.e., a parametric quadratic program defined metric, is used to formulate a nonsmooth CBF condition based on the directional derivative of the KKT solution. How to extend it to other problems is also not obvious. 3) Recently, the nonsmoothness in PCA has been avoided by circumscribing polytopes into strictly convex shapes to construct differentiable optimization defined CBFs [11, 12, 13] or by proposing some nonconservative approximation of signed distance field to establish the optimization-free smooth CBF [14]. This indirect approach avoids the challenges of nonsmooth parametric optimization, transforming the original simple LP into a nonlinear program. In summary, the existing works about parametric optimization defined CBF methods [10, 11, 12, 13, 14, 15, 16] could not be directly applied to address the difficulties of nonsmoothness in FSM problems.
@@
This paper extends the study of Feasible Space Monitoring [9] and addresses the unresolved difficulties in nonsmooth parametric optimization based CBFs [10, 11, 12, 13, 14, 15, 16]. First, a broader class of control problems is formulated, called Polytope Volume Monitoring (PVM). By identifying this kind of problems, FSM problem [9] and Polytopic Intersectability Maintenance (PIM), i.e., the opposite problem of PCA [11, 12, 13, 14, 15], can be regarded as different special examples of PVM, as seen in Fig. 1. This generalization enhances the applicability of the proposed method to a wider range of problems. Second, to address difficulties in analysis and synthesis caused by inherent nonsmoothness, this paper attempts to establish a bridge between the directional derivative defined nonsmooth CBF [17] and parametric optimization theory [18]. In detail, our main contributions can be summarized as follows:
@@
    Within this unified framework, a parametric LP based nonsmooth barrier function candidate is established for PVM, based on which sufficient conditions of being a nonsmooth CBF are proposed. Then a slacked QP-based safety filter with guaranteed feasibility is proposed to solve the PVM problem. The above nonsmooth analysis and considerations are not addressed in [9].
@@
The paper is organized as follows. The CBF method is briefly reviewed in Section II, and problem formulation and preliminary analysis are given in Section III. Section IV reports the main results. Numerical simulation is performed in Section V. Appendix includes the referred nonsmooth parametric optimization theory and proofs.
Notations: Throughout this paper, $\operatorname{col}(x_{1},x_{2}):=[x_{1}^{\top},x_{2}^{\top}]^{\top}$ for any two column vectors $x_{1},x_{2}$. For any vector $x$, $\|x\|$ denotes its Euclidean norm. $|\mathcal{J}|$ denotes the cardinality of any finite number set $\mathcal{J}$. 00 and $\emptyset$ denote the zero vector and empty set respectively in suitable dimension. The interior and boundary of a set $\mathcal{C}$ are indicated by $\operatorname{Int}(\mathcal{C})$ and $\partial\mathcal{C}$. A continuous function $\alpha:\mathbb{R}\rightarrow\mathbb{R}$ is called in extended class-$\mathcal{K}$ if it is strictly increasing and with $\alpha(0)=0$ [1]. For any function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$, if exists, its generalized gradient at $x$ is denoted by $\partial f(x)$, and $f^{\prime}(x;d)$ denotes its directional derivative at $x$ along direction $d$. For brevity, its partial derivative $\frac{\partial f}{\partial x}(x)$, expressed as a row vector, is denoted by $\partial_{x}f$ sometimes.
@@ II Preliminaries: Control Barrier Function
Given a nonempty closed set $\mathcal{C}\subset\mathbb{R}^{n}$ with no isolated point, a locally Lipschitz continuous function $h$ : $\mathbb{R}^{n}\rightarrow\mathbb{R}$ is called a NBF candidate on set $\mathcal{C}$ if
@@
Function $h$ is called a differentiable CBF to (1) if BF candidate $h$ is further differentiable and there exists an enxtended $\mathcal{K}$-class function $\alpha$ such that for all $x$ it holds $\sup_{u\in\mathcal{D}_{u}}\left\{\frac{\partial h}{\partial x}(x)f_{0}(x,u)\right\}\geq-\alpha(h(x)).$ For the demands of this paper, we only require from now on that $h$ is Lipschitz continuous and the (one-sided) directional derivative of $h(x)$ at point $x$ along the direction $\dot{x}$ exists, i.e., the following limit is finite and abbreviated as
$h^{\prime}(x(t);\dot{x}(t))$ $:=\lim_{\sigma\rightarrow 0^{+}}\frac{h(x(t)+\sigma\dot{x}(t))-h(x(t))}{\sigma}.$ (3)
@@
A locally Lipschitz continuous function $h:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is a nonsmooth control barrier function (NCBF) for system (1) if $h$ is a NBF candidate on the closed set $\mathcal{C}\subset\mathbb{R}^{n}$, and there exists an extended class-$\mathcal{K}$ function $\alpha:\mathbb{R}\rightarrow\mathbb{R}$, an open set $\mathcal{D}$ that satisfies $\mathcal{C}\subset\mathcal{D}\subseteq\mathbb{R}^{n}$, such that:
@@
In [17], the same limitation value as $h^{\prime}(x;f_{0}(x,u))$ is used to define NCBF, yet without introducing directional derivative formally. In this paper, the directional derivative defined NCBF (4) is introduced to build a bridge between the NCBF theory and the parametric optimization theory.
@@
[17, Theorem 1]  If $h$ is a NCBF for system (1), then the closed set $\mathcal{C}$ is forward control-invariant, that is, there exists $u(t)\in\mathcal{D}_{u}$, such that $x_{0}\in\mathcal{C}$ implies that $x(t)\in\mathcal{C},~{}\forall t\in\left[0,T\right]$, for every solution.
@@
where $u_{0}$ is any given continuous nominal control law.
When $h(x)$ is no longer differentiable but directional derivative $h^{\prime}(x;f_{0}(x,u))$ exists, the inequality (5b) may no longer be a linear constraint of input $u$.
@@
III Problem Formulation and Analysis
@@ III-A Problem Statement
$y$ $=\Phi(x):=\left\{c\in\mathbb{R}^{l}\mid a_{j}(x)^{\top}c\leq b_{j}(x),j\in\mathcal{J}\right\},$ (7b)
$u$ $\in\Psi(x):=\left\{u\in\mathbb{R}^{m}\mid\mathfrak{a}_{\mathfrak{i}}(x)^{\top}u\leq\mathfrak{b}_{\mathfrak{i}}(x),\mathfrak{i}\in\mathfrak{J}\right\},$ (7c)
where $\mathcal{J}:=\{1,2,...,N\}$ and $\mathfrak{J}:=\{1,2,...,\mathfrak{N}\}$. For convenience, the functions appearing in definitions of $\Phi(x)$ and $\Psi(x)$ are assumed to be continuously differentiable, and the following assumptions are imposed to develop an implementable controller.
@@
The state-dependent polytopic sets $\Phi(x)$, $\Psi(x)$ are both nonempty and uniformly bounded for all $x$, i.e., there exist compact sets $\mathcal{D}_{y}\subset\mathbb{R}^{l}$ and $\mathcal{D}_{u}\subset\mathbb{R}^{m}$, such that $\Phi(x)\subseteq\mathcal{D}_{y}$ and $\Psi(x)\subseteq\mathcal{D}_{u}$ respectively. Additionally, vectors $a_{j}(x),j\in\mathcal{J}$ are nonzero at any $x$.
Without loss of generality, Assumption 1 is made for some technical considerations. Especially, the non-emptiness of $\Phi(x)$ is essential to guarantee that some metric is well-defined in the following sections. In fact, this paper will attempt to synthesize a controller to prevent the interior of $\Phi(x(t))$ from vanishing, as formally stated below.
Figure 1: Illustration of the PVM problem and its examples. (a) Problem statement and preliminary analysis. (b) Example 1 in input space. (c) Example 2 in state space.
Polytope Volume Monitoring Problem: Consider the system (7) with a polytopic set-valued output $\Phi(x)$ and a polytopic input constraint set $\Psi(x)$. The goal is to design a control law for $u\in\Psi(x)$ to prevent the volume of the polytope $\Phi(x(t))$ from being less than a pre-defined positive value, i.e.,
$\mathcal{V}(x(t)):=\int_{\Phi(x(t))}\mathrm{d}s\geq\epsilon,~{}\forall t\geq 0,$ (8)
where $\epsilon>0$ is any user-defined threshold value.
There are some typical situations where the volume of state-dependant polytope $\Phi(x(t))$ might be zero if the input $u$ is not designed properly.
@@
In Fig. 1(b), given the polytopic input constraint set $\mathcal{U}$, by constructing the set composed of multiple differentiable CBFs $h_{i},i\in\mathcal{I}$: $\mathcal{U}_{c}(x):=\{u\mid\frac{\partial h_{i}}{\partial x}(x)\left(f(x)+g(x)u\right)\geq-\alpha_{i}(h_{i}(x)),i\in\mathcal{I}\}$, the FSM problem aims to design $u$ such that
$\mathcal{U}\cap\mathcal{U}_{c}(x(t))\neq\emptyset,~{}\forall t.$ (9)
By choosing $\Phi(x)=\Psi(x)=\mathcal{U}\cap\mathcal{U}_{c}(x)$, feasible space monitoring of any QP-based controllers can be regarded as case studies of the formulated PVM problem.
@@
Consider two standard nonholonomic vehicles $i,j$, for each $\mathfrak{i}\in\{i,j\}$, $x^{\mathfrak{i}}=[{p}_{x}^{\mathfrak{i}},{p}_{y}^{\mathfrak{i}},{\theta}^{\mathfrak{i}}]^{\top}$ and $u^{\mathfrak{i}}=[v^{\mathfrak{i}},\omega^{\mathfrak{i}}]^{\top}\in\mathcal{U}^{\mathfrak{i}}$ are state and input vectors respectively, where $\mathcal{U}^{\mathfrak{i}}$ denotes a hyperrectangle. Let sets $\mathcal{P}^{i}(x^{i})$ and $\mathcal{P}^{j}(x^{j})$ represent their physical domain or sensor ranges respectively, as shown in Fig. 1(c). The goal of polytopic intersectability maintenance is to maintain the intersection of two polytopes nonempty:
$\mathcal{P}^{i}(x^{i}(t))\cap\mathcal{P}^{j}(x^{j}(t))\neq\emptyset,~{}\forall t,$ (10)

### v3

@@ Abstract
Motivated by the latest research on feasible space monitoring of multiple control barrier functions (CBFs) as well as polytopic collision avoidance, this paper studies the Polytope Volume Monitoring (PVM) problem, whose goal is to design a control law for inputs of nonlinear systems to prevent the volume of some state-dependent polytope from decreasing to zero. Recent studies have explored the idea of applying Chebyshev ball method in optimization theory to solve the case study of PVM; however, the underlying difficulties caused by nonsmoothness have not been addressed. This paper continues the study on this topic, where our main contribution is to establish the relationship between nonsmooth CBF and parametric optimization theory through directional derivatives for the first time, to solve PVM problems more conveniently. In detail, inspired by Chebyshev ball approach, a parametric linear program (PLP) based nonsmooth barrier function candidate is established for PVM, and then, sufficient conditions for it to be a nonsmooth CBF are proposed, based on which a quadratic program (QP) based safety filter with guaranteed feasibility is proposed to address PVM problems. Finally, a numerical simulation example is given to show the efficiency of the proposed safety filter.
@@ I Introduction
The appearance of multiple CBFs and input constraints would impose great challenges in the feasibility of the resulting QP-based control laws, which has been paid much attention recently [5, 6, 7, 8, 9]. Some primary works study sufficient conditions of the compatibility of multiple CBFs [6, 7]. More recently, in [8], the nonsmooth CBF, dating backing to [5], is introduced as a systematic tool to address challenges in guaranteeing the feasibility and continuity. Meanwhile, a simpler way of addressing the feasibility problem appears in [9]. In detail, “Feasible CBF Space” is defined to describe the state-dependant polytope formulated by CBF conditions and input constraints. Then, “Feasible Space Monitoring” (FSM) problem is established to monitor the feasibility of the resulting QP-based control laws, for which Chebyshev ball and inscribing ellipsoid methods are proposed to build parametric convex program based CBFs. For engineering implementations, the differentiable optimization toolbox is applied to approximate the gradient of the resulting (nonsmooth implicit) barrier function candidate by some heuristic gradient values. As concluded in [9], the nonsmooth control theoretic considerations are left for future work. To the best of our knowledge, this problem has been open until now. Such observation motivates this paper.
@@
Meanwhile, researchers have been gradually interested in building CBFs via parametric optimization-defined metrics for other problems, such as network connectivity maintenance [10] and polytopic collision avoidance (PCA) [11, 12, 13, 14, 15, 16]. However, it is difficult to extend these results to solve the above open problem in FSM. Roughly speaking, there are three potential ways, yet all meet significant difficulties: 1) In [10], the Laplacian’s eigenvalue is chosen as a barrier function candidate, then, a nonsmooth CBF condition in the generalized gradient form is formulated. The idea of building nonsmooth CBF conditions based on generalized gradient can only be applied to specific problems, and it is not applicable to other parametric optimization problems. 2) In [15, 16], the minimum distance field for PCA, i.e., a parametric quadratic program defined metric, is used to formulate a nonsmooth CBF condition based on the directional derivative of the KKT solution. How to extend it to other problems is also not obvious. 3) Recently, the nonsmoothness in PCA has been avoided by circumscribing polytopes into strictly convex shapes to construct differentiable optimization defined CBFs [11, 12, 13] or by proposing some nonconservative approximation of signed distance field to establish the optimization-free smooth CBF [14]. This indirect approach avoids the challenges of nonsmooth parametric optimization, transforming the original simple LP into a nonlinear program. In summary, the existing works about parametric optimization defined CBF methods [10, 11, 12, 13, 14, 15, 16] could not be directly applied to address the difficulties of nonsmoothness in FSM problems.
@@
This paper extends the study of Feasible Space Monitoring [9] and addresses the unresolved difficulties in nonsmooth parametric optimization based CBFs [10, 11, 12, 13, 14, 15, 16]. First, a broader class of control problems is formulated, called Polytope Volume Monitoring (PVM). By identifying this kind of problems, FSM problem [9] and Polytopic Intersectability Maintenance, i.e., the opposite problem of PCA [11, 12, 13, 14, 15], can be regarded as different special examples of PVM. This generalization enhances the applicability of the proposed method to a wider range of problems. Second, to address difficulties in analysis and synthesis caused by inherent nonsmoothness, this paper attempts to establish a bridge between the directional-derivative-defined nonsmooth CBF [17] and parametric optimization theory [18]. In detail, our main contributions can be summarized as follows:
@@
    Within this unified framework, a parametric linear program based nonsmooth barrier function candidate is established for PVM, based on which sufficient conditions of being a nonsmooth CBF are proposed. Then a a feasibility-guaranteed QP-based safety filter is proposed to solve the PVM problem. The above nonsmooth analysis and considerations are not addressed in [9].
@@
The paper is organized as follows. The CBF method is briefly reviewed in Section II, and problem formulation and preliminary analysis are given in Section III. Section IV reports the main results. The referred nonsmooth parametric optimization theory is included in Appendix.
Notations: Throughout this paper, $\operatorname{col}(x_{1},x_{2})\mathrel{\mathop{\ordinarycolon}}=[x_{1}^{\top},x_{2}^{\top}]^{\top}$ for any two column vectors $x_{1},x_{2}$. For any vector $x$, $\|x\|$ denotes its Euclidean norm. $|\mathcal{J}|$ denotes the cardinality of any finite number set $\mathcal{J}$. 0 and $\emptyset$ denote the zero vector and empty set respectively in suitable dimension. The interior and boundary of a set $\mathcal{C}$ are indicated by $\operatorname{Int}(\mathcal{C})$ and $\partial\mathcal{C}$. A continuous function $\alpha\mathrel{\mathop{\ordinarycolon}}\mathbb{R}\rightarrow\mathbb{R}$ is called in extended class-$\mathcal{K}$ if it is strictly increasing and with $\alpha(0)=0$ [1]. For any function $f\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{n}\rightarrow\mathbb{R}$, if exists, its generalized gradient at $x$ is denoted by $\partial f(x)$, and $f^{\prime}(x;d)$ denotes its directional derivative at $x$ along direction $d$. For brevity, its partial derivative $\frac{\partial f}{\partial x}(x)$, expressed as a row vector, is denoted by $\partial_{x}f$ sometimes.
@@ II Preliminaries: Control Barrier Function
Given a nonempty closed set $\mathcal{C}\subset\mathbb{R}^{n}$ with no isolated point, a locally Lipschitz continuous function $h(\cdot)\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{n}\rightarrow\mathbb{R}$ is called a NBF candidate on set $\mathcal{C}$ if
@@
For the demands of this paper, we only require from now on that $h$ is locally Lipschitz continuous and the (one-sided) directional derivative of $h(x)$ at point $x$ along any direction $d$ exists, i.e., the following limit is finite and abbreviated as
$h^{\prime}(x;d)\mathrel{\mathop{\ordinarycolon}}=\lim_{\sigma\rightarrow 0^{+}}\frac{h(x+\sigma d)-h(x)}{\sigma}.$
@@
A locally Lipschitz continuous function $h\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{n}\rightarrow\mathbb{R}$ is a nonsmooth control barrier function (NCBF) for system (1) if $h$ is a NBF candidate on the closed set $\mathcal{C}\subset\mathbb{R}^{n}$, and there exists an extended class-$\mathcal{K}$ function $\alpha\mathrel{\mathop{\ordinarycolon}}\mathbb{R}\rightarrow\mathbb{R}$, an open set $\mathcal{D}$ that satisfies $\mathcal{C}\subset\mathcal{D}\subseteq\mathbb{R}^{n}$, such that:
@@
In [17], the same limitation value as $h^{\prime}(x;f_{0}(x,u))$ is used to define NCBF, yet without introducing the concept of directional derivative. In this paper, the directional derivative defined NCBF (3) is introduced to build a bridge between the NCBF theory and the parametric optimization theory.
@@
[17, Theorem 1]  If $h$ is a NCBF for system (1), then the closed set $\mathcal{C}$ is forward control-invariant, that is, there exists a measurable and locally bounded function $u(t)\in\mathcal{D}_{u}$, such that $x_{0}\in\mathcal{C}$ implies that $x(t)\in\mathcal{C},~\forall t\in\left[0,T\right]$ for every solution.
@@
where $u_{0}$ is any given continuous nominal control law. When $h(x)$ is no longer differentiable but the directional derivative $h^{\prime}(x;f_{0}(x,u))$ exists, the inequality (4b) may no longer be a linear constraint of input $u$.
@@
(a) Polytope volume monitoring problem and examples.
(b) The structure of the proposed safety filter.
Figure 1: A summary of the polytope volume monitoring problem and the parametric linear program based control barrier function.
III Problem Formulation
@@ III-A Problem Statement
$y$ $=\Phi(x)\mathrel{\mathop{\ordinarycolon}}=\left\{c\in\mathbb{R}^{l}\mid a_{j}^{\top}(x)c\leq b_{j}(x),j\in\mathcal{J}\right\},$ (6b)
$u$ $\in\Psi(x)\mathrel{\mathop{\ordinarycolon}}=\left\{u\in\mathbb{R}^{m}\mid\mathfrak{a}_{\mathfrak{i}}^{\top}(x)u\leq\mathfrak{b}_{\mathfrak{i}}(x),\mathfrak{i}\in\mathfrak{J}\right\},$ (6c)
where $\mathcal{J}\mathrel{\mathop{\ordinarycolon}}=\{1,2,...,N\}$ and $\mathfrak{J}\mathrel{\mathop{\ordinarycolon}}=\{1,2,...,\mathfrak{N}\}$. For convenience, the functions appearing in definitions of $\Phi(x)$ and $\Psi(x)$ are assumed to be continuously differentiable, and without loss of generality, the following assumptions are imposed to develop an implementable controller.
@@
The polytopic input constraint set $\Psi(x)$ is always nonempty and the vectors $a_{j}(x),j\in\mathcal{J}$ are always nonzero. The state-dependent polytopic set $\Phi(x)$ is uniformly bounded, i.e., there exists a compact set $\mathcal{D}_{y}\subset\mathbb{R}^{l}$, such that $\Phi(x)\subseteq\mathcal{D}_{y},\forall x\in\mathbb{R}^{n}$.
It is worth noting that any emptyset is naturally bounded, thus it is not required in Assumption 1 that $\Phi(x)$ is always nonempty. Instead, we allow there exist some points $x$ at which $\Phi(x)=\emptyset$. In fact, this paper will attempt to synthesize a controller to prevent the interior of $\Phi(x(t))$ from vanishing, as formally stated below.
Polytope Volume Monitoring Problem: Consider the system (6) with a polytopic set-valued output $\Phi(x)$ and a polytopic input constraint set $\Psi(x)$. The goal is to design a control law for $u$ to prevent the volume of $\Phi(x(t))$ from decreasing to zero, i.e., to ensure the nonemptness of the interior of $\Phi(x(t))$:
$x(t)\in\mathcal{A}\mathrel{\mathop{\ordinarycolon}}=\{x\mid\operatorname{Int}(\Phi(x))\neq\emptyset\},\quad\forall t\geq 0.$ (7)
There are some typical situations where the state-dependant polytope $\Phi(x(t))$ might be empty if the input $u$ is designed improperly.
@@
In Fig. 1 (a), given the polytopic input constraint set $\mathcal{U}$, by constructing the set composed of multiple differentiable CBFs $h_{i},i\in\mathcal{I}$: $\mathcal{U}_{c}(x)\mathrel{\mathop{\ordinarycolon}}=\{u\mid\frac{\partial h_{i}}{\partial x}(x)\left(f(x)+g(x)u\right)\geq-\alpha_{i}(h_{i}(x)),i\in\mathcal{I}\}$, the FSM problem aims to design $u$ such that
$\mathcal{U}\cap\mathcal{U}_{c}(x(t))\neq\emptyset,~\forall t\geq 0.$ (8)
By choosing $\Phi(x)=\Psi(x)=\mathcal{U}\cap\mathcal{U}_{c}(x)$, FSM of any QP-based controllers can be regarded as the case studies of the formulated PVM problem.
@@
Consider two standard nonholonomic vehicles $i,j$, for each $\mathfrak{i}\in\{i,j\}$, $x^{\mathfrak{i}}=[{p}_{x}^{\mathfrak{i}},{p}_{y}^{\mathfrak{i}},{\theta}^{\mathfrak{i}}]^{\top}$ and $u^{\mathfrak{i}}=[v^{\mathfrak{i}},\omega^{\mathfrak{i}}]^{\top}\in\mathcal{U}^{\mathfrak{i}}$ are state and input vectors respectively, where $\mathcal{U}^{\mathfrak{i}}$ denotes a hyperrectangle. Let $\mathcal{P}^{i}(x^{i}),\mathcal{P}^{j}(x^{j})$ denote their physical domain or sensor ranges, as in Fig. 1 (a). The goal of PIM is to maintain the intersection of two polytopes nonempty:
$\mathcal{P}^{i}(x^{i}(t))\cap\mathcal{P}^{j}(x^{j}(t))\neq\emptyset,~\forall t\geq 0,$ (9)